Chuck owns a sandwich shop that is open Monday through Saturday. He is thinking of opening the shop for business on Sundays as well, in which case he will make an additional $4,000 per month. . He actually owns the shop. If he rented it, rather than using it as a sandwich shop, he would make a rental income of S5,000 per month. .He pays $1,500 per month to a bank, regardless of the number of days the shop is open for business, for the loan he took out a while back to buy a new commercial refrigerator Two years ago he paid a one-time fee of $2,000 to renew his business license. .Four years ago he paid $7,000 to paint the shop and repair the floor .The operating cost (labor, electricity, water, materials) will increase by S3,000 per month if he opens the shop for business on Sundays. .He will lose $1,700 per month that he currently makes teaching business administration on Sundays at a local school. Assume that these are the only costs and benefits to consider Question 10 Consider the information above. With respect to Chucks decision, what is the sunk cost? A. $7,000 B. $7,800 C. $8,000 D. $9,000 E. $9,500 Question 11 Consider the information above. With respect to Chucks decision, what is the total fixed cost? A. $2,700 B. $4,500 C. $5,000 D. $5,700 E. $6,500 Question 12 Consider the information above. With respect to Chucks decision, what is the total marginal cost? A. $3,000 B. $3,700 C. $4,700 D. $3,900 E. $5,000 Question 13 Consider the information above. With respect to Chucks decision, what is the net benefit of opening the shop for business on Sundays (net benefit benefit minus cost)? A. $1,000 B. - $700 C. +$700 D. +$800 E. +$1,000 Page 4 of 10

Please solve and show/explain work for questions #10- 13 for a rating. The answers are 10. D , 11. E, 12. C, and 13. B
Answer #1

10. Sunk cost is the cost that has already been incurred and cannot be recovered.

Sunk cost = $(2000+7000)= $9000. Hence, option(D) is correct.

11. Total fixed cost is the cost that has been incurred when there is no production.

Total fixed cost = $(5000+1500)= $6500. Hence, option(E) is correct.

12. Marginal cost is the additional cost that has been incurred in the production of additional quantity.

Total marginal cost =$(3000+1700) = $4700. Hence, option(C) is correct.

13. Net benefit of opening the shop for business on Sundays = Benefit - Cost

Benefit = $4000 (given)

Cost = $4700

Net benefit = $(4000-4700)= -$700. Hence, option(B) is correct.